# Stability of quasi-particle creation and multiband geometry in fractional Chern insulators under magnetic fields

### v1

@@ Abstract
We study creation of quasi-particles in fractional Chern insulators (FCI) under magnetic fields. We consider two representative models, the Kapit-Mueller model and the checkerboard model, which have distinct band properties in terms of the quantum geometry. The former satisfies the so-called ideal condition and well mimics the lowest Landau level, while the latter is not ideal for realization of FCI states. It is found within exact diagonalization that both quasi-holes and quasi-electrons are stably created by the magnetic fields in the Kapit-Mueller model. On the other hand, stability of the quasi-particle creation depends on directions of the magnetic field in the checkerboard model. Although the quasi-electron creation is stable under a magnetic field, the quasi-hole creation and the underlying FCI state are unstable for the opposite field direction, leading to a field-induced non-FCI state. We point out that this difference can be understood based on the multiband quantum geometry in the presence of the magnetic fields.
@@ I Introduction
The FCIs have been observed experimentally in moire materials [28, 29, 30, 31, 32, 33, 34]. Especially for MoTe2 and rombohedral pentalayer graphene, the fractional anomalous quantum Hall effects at exactly zero fields and stable plateaus of the Hall resistance in magnetic fields have been reported. Although the applied magnetic fields themselves are small, the corresponding fluxes per moire unit cell can be large because the moire periods are much longer than typical lattice constants. Therefore, even small magnetic fields may have non-trivial effects on FCIs, especially on the quantum Hall plateaus. The existence of a plateau of the quantum Hall conductance is one of the most characteristic and fundamental properties in FQHE. The standard understanding of this phenomenon is that quasi-particles with fractional charges are created when the magnetic field is varied in an FQH state, but they are spatially localized by disorder and consequently the quantized Hall effect robustly remains [35, 36]. Similar mechanisms would work also in FCIs in magnetic fields. However, there are various systems for FCIs which are distinct in terms of the quantum geometry and vortexability. It is not trivial how such differences could influence stability of the quantized Hall effect in FCIs under magnetic fields. Especially, the magnetic field can affect the underlying band structure and, consequently the quantum geometry could change possibly leading to instability of FCIs in some systems. This is in sharp contrast to the creation of quasi-particles by changing the electron filling of FCIs in the absence of magnetic fields, where the underlying band structures are kept intact [1, 37, 38, 39].
In this study, we investigate creation of quasi-particles and stability of FCI states under magnetic fields in two representative models which have distinct band properties, namely the Kapit-Mueller model and the checkerboard model [23, 40]. The former model satisfies the ideal condition, while the latter does not. We demonstrate by the exact diagonalization that the two models show different behaviors under magnetic fields. Both of the quasi-holes and quasi-electrons are stable in the Kapit-Mueller model, while the quasi-holes are unstable in the checkerboard model. The difference in the stability can be attributed to the quantum geometries of multiband structures under magnetic fields.
@@ II.1 Brief overview of quantum geometry
where $\alpha,\beta\in\{x,y\}$ and $\bm{k}$ is a wave vector in the Brillouin zone. $A$ is the area of the Brillouin zone. |$\ket{u_{\bm{k}}}$ is the cell-periodic function for the band which we focus on, and ${\mathcal{P}}(\bm{k})=\ket{u_{\bm{k}}}\bra{u_{\bm{k}}}$ | is the corresponding projection operator. $g(\bm{k})$ and $\mathcal{F}(\bm{k})$ are Fubini-Study metric and Berry curvature, respectively. They are given as real and imaginary parts of the quantum geometric tensor, and known to satisfy the following inequality,
@@ II.2 Two lattice models
We introduce two models to understand the quasi-particle creation and the stability of the FCIs against magnetic fields: the Kapit-Mueller model [23, 24] and the checkerboard model [40]. Both models have flat bands with the Chern number $|C_{\text{s}}|=|(1/2\pi A)\int d^{2}k\mathcal{F}_{xy}|=1$ and realize the FCIs in the absence of magnetic fields, but they have different quantum geometries. In presence of interactions, FCI states are driven in these models. In this study, we consider spinless fermions with the nearest neighbor interaction
@@
where $z_{k}=x_{k}+iy_{k}$ is the lattice complex coordinate for the site $k$ and $z=z_{j}-z_{k}=x+iy$. $\phi_{0}$ is a model parameter which is distinguished from the applied flux and chosen to be $\phi_{0}=1/3$. For this parameter, there are three energy bands in the absence of the magnetic field, and the lowest energy band is exactly flat and has a Chern number $C_{s}=1$. In addition, it satisfies the ideal condition and thus the single-particle Bloch function of this band is a holomorphic function. $\phi$ is the applied flux per plaquette, and there is no distinction between the flux and the magnetic field in this study since the lattice constant is the length unit. Note that, since $\phi$ is not included in the Gaussian factor in the hopping integral $J(z_{j},z_{k})$, the Poisson summation rule [23, 41] for hopping integrals does not hold at $\phi\neq 0$ and the ideal condition is not necessarily satisfied under the magnetic field. Nevertheless, deviation from the ideal condition is small as will be discussed later in Sec. V. $G_{jk}$ is the phase factor corresponding to the magnetic field $\phi$ in addition to the model parameter $\phi_{0}=1/3$. We use the string gauge [42, 43, 44, 45],
@@
(a) Kapit-Mueller model Refer to caption
(b) Checkerboard model Refer to caption
@@ III Summary of results
Before presenting detailed calculation results, we first summarize the key results here. Quasi-particles can be created by the magnetic field $\phi$, and they can be localized by an additional pinning potential $V_{p}$ corresponding to disorder. We show expected phase diagrams of the two models with the pinning potential $V_{p}\neq 0$ in Fig. 2 based on the exact diagonalization and the quantum geometry under magnetic fields. Generally, a magnetic field increases or decreases the number of (nearly degenerate) single-particle states $N_{\text{orb}}$ according to the Streda’s formula [48, 49, 50], $N_{\text{orb}}\to N_{\text{orb}}+C_{s}N_{\phi}$, where $C_{\text{s}}$ is the Chern number of the band considered and $N_{\phi}$ is the number of the added flux for the entire system. If $N_{\text{orb}}$ increases (decreases) under a magnetic field, the filling $\nu=N/N_{\text{orb}}$ decreases (increases), which can lead to quasi-hole (quasi-electron) creation. In the Kapit-Mueller model, since the Chern number of the lowest energy band is $+1$, adding a positive magnetic field induces quasi-hole creation, while a negative magnetic field leads to quasi-electron creation. In the checkerboard model, the Chern number of the lowest energy band is $-1$, and a positive (negative) magnetic field corresponds to quasi-electron (quasi-hole) creation. The quasi-holes can be localized by a pinning potential $V_{p}>0$ while the quasi-electrons are by $V_{p}<0$ in our notation. Within the exact diagonalization for limited system sizes in the present study, the number of the created quasi-particles is $N_{\text{qp}}=1$ in each model. In a thermodynamically large system, the number of the flux quanta is $|N_{\phi}|\gg 1$ for a fixed magnetic field $\phi$ and there will be $|N_{\phi}|$ quasi-particles which are localized by a pinning potential. In Fig. 2, we have assumed that the quasi-particles are well localized and effects of interactions between the pinned quasi-particles are negligible.
As seen in Fig. 2(a) for the Kapit-Mueller model, the FCI state is stable under the magnetic field irrespective of its direction. More precisely, except for very small interactions, quasi-particles are stably created by the magnetic field and then localized by the pinning potential. Consequently, the fractionally quantized Hall conductance is maintained over a wide range of parameters, which means that the FCI phase is stable. The phase diagram is almost symmetric for $+\phi$ and $-\phi$, and thus the creations of quasi-holes ($\phi>0$) and quasi-electrons ($\phi<0$) are both stable. This is a reasonable result, since the Kapit-Mueller model is a lattice model which well mimics the lowest Landau level in the continuum system and basic properties are similar to the corresponding FQHE.
On the other hand, in the checkerboard model, the quasi-electron creation under $\phi>0$ is stable except for small interactions, while the quasi-hole creation for $\phi<0$ is unstable even for moderately strong interactions. In the latter case, the FCI state is destabilized by the magnetic field and a non-FCI state is realized. The non-FCI state at very small interactions is smoothly connected to the non-interacting Fermi liquid at $U=0$. However, it is difficult to fully clarify nature of the non-FCI state within the present calculations, and exact identification of the non-FCI state is left for a future study. We stress that the asymmetry with respect to ±$\pm\phi$ is a characteristic feature of a lattice system which is absent in the standard FQHE in the continuum model.
@@
Figure 2: Expected phase diagrams for (a) the Kapit-Mueller model and (b) the checkerboard model with the pinning potential $V_{p}\neq 0$. $\phi$ is the applied magnetic field and $U$ is the strength of the nearest-neighbor interaction. The strength of the pinning potential is chosen so that the created quasi-particles are well localized. It is assumed that effects of interactions between the localized quasi-particles are negligible in a thermodynamically large system.
@@ IV.1 Characterization of FCI
In this section, we study the FCI with use of the exact diagonalization. First, we briefly explain how to characterize an FCI state within the exact diagonalization. To examine stability of the FCI states against magnetic fields, we apply one flux quantum $N_{\phi}=\pm 1$ uniformly over the entire system and discuss quasi-particles induced by the magnetic field based on the “counting-rule” on the energy spectra [1, 51, 52]. Previous works examined quasi-hole and quasi-electron creations in FCIs at $\phi=0$ based on the counting-rule, when the number of unit cells is increased or decreased [37, 38, 39]. When the system size is changed with a fixed number $N$ of the fermions, the number of the nearly degenerate single-particle states $N_{\text{orb}}$ and the filling $\nu=N/N_{\text{orb}}$ varies, while the single-particle band structure itself is unchanged. Instead of varying the system size, a change in the number of the single-particle orbits can be induced by the magnetic field according to the Streda’s formula, $N_{\text{orb}}\to N_{\text{orb}}+C_{s}N_{\phi}$. In contrast to the system-size induced change of $N_{\text{orb}}$, the system does not keep the (magnetic) translation symmetry under the additional one flux quantum $N_{\phi}=\pm 1$ [44]. However, one can still regard the nearly degenerate single-particle orbits as a nearly flat band and discuss the stability of the FCI in a manner similar to those in the previous studies at $\phi=0$.
Since the single-particle Chern number is $|C_{s}|=1$ in the present models, a single flux quantum leads to the change $N_{\text{orb}}\to N_{\text{orb}}\pm 1$. This gives rise to creation of one quasi-hole or quasi-electron with the fractional charge $|e^{*}|=1/3$. In case of the quasi-hole creation, the ground state degeneracy #$\#_{\text{qh}}$ under the magnetic field obeys the counting rule given by  [1]
@@
where $N_{\text{orb}}$ is the number of the low energy single-particle states in the presence of the magnetic field. For quasi-electron creation, no simple formula for the ground state degeneracy #$\#_{\text{qe}}$ is known, but we can verify the counting rule of the quasi-electrons by referring to energy spectra of the corresponding FQHE states. More precisely, we calculate #$\#_{\text{qe}}$ for the LLL in the electron gas model and also for fermions in the Kapit-Mueller model with the Poisson summation rule which is a lattice realization of the LLL [23, 41]. We numerically find that #$\#_{\text{qe}}=11$ for $(N=4,N_{\text{orb}}=11)$, #$\#_{\text{qe}}=14$ for $(N=5,N_{\text{orb}}=14)$, and #$\#_{\text{qe}}=17$ for $(N=6,N_{\text{orb}}=17)$.
The quasi-particles created by the uniform magnetic field are extended over the entire system. They can be spatially localized in the presence of an additional pinning potential $V_{p}$ which couples with the quasi-particle charge [35]. To this end, we first perform the exact diagonalization to obtain low energy states |$\ket{\Psi_{1}},\ket{\Psi_{2}},\cdots,\ket{\Psi_{N_{\text{cut}}}}$ without the pinning potential [37, 38], where $N_{\text{cut}}=\#_{\text{qh}}$ or #$\#_{\text{qe}}$ is a cut-off for the low energy states. Then, we take $V_{p}$ into account within this low energy sector by diagonalizing the projected Hamiltonian with the pinning potential,
@@
where $H$ is either $H_{\text{KM}}$ or $H_{\text{CB}}$. $P$ is the projection onto the low energy states {|$\{\ket{\Psi_{n}}\}_{n=1}^{N_{\text{cut}}}$ and $j_{0}$ is the position of the pinning potential. The projection enables us to efficiently focus on the low energy sector and also greatly reduces computational costs for the analysis of the pinning potential [37, 38]. When the quasi-particle created by $\phi\neq 0$ is localized by the pinning potential, it does not contribute to the Hall conductance. Correspondingly, the ground state degeneracy of $\tilde{H}$ will be three fold for the present $\nu=1/3$ system. In this sense, the ground states of $\tilde{H}$ belong to the same FCI phase of $H$ at $\phi=0$. However, the quasi-particle creation can be unstable and a magnetic field may destabilize the underlying FCI state. In this case, the counting rule does not hold and the ground states of $\tilde{H}$ are no longer FCIs.
@@
The variance is exponentially small if the ground states are FCIs, while it can be large for gapless states. In this way, we can identify FCI states within the exact diagonalization for small system sizes. It turns out that, even if the ground state of $\tilde{H}$ is an FCI, $\mathcal{V}[\mathcal{C}]$ diverges when we use $N_{\text{cut}}<\#_{\text{qh}},\#_{\text{qe}}$. When $N_{\text{cut}}=\#_{\text{qh}},\#_{\text{qe}}$, the variance $\mathcal{V}[\mathcal{C}]$ is suppressed in the FCI phase.
In this study, we consider the systems with the number of the fermions $N=4\sim 6$. This is relatively small compared to those in the previous studies where quasi-particles are created by the increase of the system sizes [1, 37, 38, 39]. Technically, one flux quantum breaks the (magnetic) translation symmetry and the size of the Hilbert space cannot be reduced by the translation symmetry [44]. However, it turns that we can discuss physical properties of the systems with $N=4\sim 6$ by carefully paying attention to finite size effects as will be discussed in the following sections. Furthermore, we have confirmed that quasi-particles are stably created by the increase of the system sizes for $N=4,5$ in consistent with the previous studies [37, 38, 39]. This implies that finite size effects are small for these system sizes.

### v2

# Stability of quasiparticle creation and multiband quantum geometry in fractional Chern insulators under magnetic fields

@@ Abstract
We study creation of quasiparticles in fractional Chern insulators (FCI) under magnetic fields. We consider two representative models, the Kapit-Mueller model and the checkerboard model, which have distinct band properties in terms of the quantum geometry. The former satisfies the so-called ideal condition and well mimics the lowest Landau level, while the latter is not ideal for realization of FCI states. It is found within exact diagonalization that both quasiholes and quasielectrons are stably created by the magnetic fields in the Kapit-Mueller model. On the other hand, stability of the quasiparticle creation depends on directions of the magnetic field in the checkerboard model. Although the quasielectron creation is stable under a magnetic field, the quasihole creation and the underlying FCI state are unstable for the opposite field direction, leading to a field-induced non-FCI state. We point out that this difference can be understood based on the multiband quantum geometry in the presence of the magnetic fields.
@@ I Introduction
The FCIs have been observed experimentally in moire materials [28, 29, 30, 31, 32, 33, 34]. Especially for MoTe2 and rombohedral pentalayer graphene, the fractional anomalous quantum Hall effects at exactly zero fields and stable plateaus of the Hall resistance in magnetic fields have been reported. Although the applied magnetic fields themselves are small, the corresponding fluxes per moire unit cell can be large because the moire periods are much longer than typical lattice constants [35, 36, 37, 38]. Therefore, even small magnetic fields may have non-trivial effects on FCIs, especially on the quantum Hall plateaus. The existence of a plateau of the quantum Hall conductance is one of the most characteristic and fundamental properties in FQHE. The standard understanding of this phenomenon is that quasiparticles with fractional charges are created when the magnetic field is varied in an FQH state, but they are spatially localized by disorder and consequently the quantized Hall effect robustly remains [39, 40]. Similar mechanisms would work also in FCIs in magnetic fields. However, there are various systems for FCIs which are distinct in terms of the quantum geometry and vortexability. It is not trivial how such differences could influence stability of the quantized Hall effect in FCIs under magnetic fields. Especially, the magnetic field can affect the underlying band structure and, consequently the quantum geometry could change possibly leading to instability of FCIs in some systems. This is in sharp contrast to the creation of quasiparticles by changing the electron filling of FCIs in the absence of magnetic fields, where the underlying band structures are kept intact [1, 41, 42, 43].
In this study, we investigate creation of quasiparticles and stability of FCI states under magnetic fields in two representative models which have distinct band properties, namely the Kapit-Mueller model and the checkerboard model [23, 44]. The former model satisfies the ideal condition, while the latter does not. We demonstrate by the exact diagonalization that the two models show different behaviors under magnetic fields. Both of the quasiholes and quasielectrons are stable in the Kapit-Mueller model, while the quasiholes are unstable in the checkerboard model. The difference in the stability can be attributed to the quantum geometries of multiband structures under magnetic fields.
@@ II.1 Brief overview of quantum geometry
where $\alpha,\beta\in\{x,y\}$ and $\bm{k}$ is a wave vector in the Brillouin zone. $A$ is the area of the Brillouin zone which has been introduced for later convenience. |$\ket{u_{\bm{k}}}$ is the cell-periodic function for the band which we focus on, and ${\mathcal{P}}(\bm{k})=\ket{u_{\bm{k}}}\bra{u_{\bm{k}}}$ | is the corresponding projection operator. $g(\bm{k})$ and $\mathcal{F}(\bm{k})$ are Fubini-Study metric and Berry curvature, respectively. They are given as real and imaginary parts of the quantum geometric tensor, and known to satisfy the following inequality,
@@ II.2 Two lattice models
(a) Kapit-Mueller model Refer to caption
(b) Checkerboard model Refer to caption
We introduce two models to understand the quasiparticle creation and the stability of the FCIs against magnetic fields: the Kapit-Mueller model [23, 24] and the checkerboard model [44]. Both models have flat bands with the Chern number $|C_{\text{s}}|=|(1/2\pi A)\int d^{2}k\mathcal{F}_{xy}|=1$ and realize the FCIs in the absence of magnetic fields, but they have different quantum geometries. In presence of interactions, FCI states are driven in these models. In this study, we consider spinless fermions with the nearest neighbor interaction
@@
where $z_{k}=x_{k}+iy_{k}$ is the lattice complex coordinate for the site $k$ and $z=z_{j}-z_{k}=x+iy$. $\phi_{0}$ is a model parameter which is distinguished from the applied flux and is chosen to be $\phi_{0}=1/3$ so that the model can efficiently describe the $\nu=1/3$-Laughlin state. For this parameter, there are three energy bands in the absence of the magnetic field, and the lowest energy band is exactly flat and has a Chern number $C_{s}=1$. In addition, it satisfies the ideal condition and thus the single-particle Bloch function of this band is a holomorphic function. $\phi$ is the applied flux per plaquette, and there is no distinction between the flux and the magnetic field in this study since the lattice constant is the length unit. Note that, since $\phi$ is not included in the Gaussian factor in the hopping integral $J(z_{j},z_{k})$, the Poisson summation rule [23, 46] for hopping integrals does not hold at $\phi\neq 0$ and the ideal condition is not necessarily satisfied under the magnetic field. Nevertheless, deviation from the ideal condition is small as will be discussed later in Sec. V. $G_{jk}$ is the phase factor corresponding to the magnetic field $\phi$ in addition to the model parameter $\phi_{0}=1/3$. We use the string gauge [47, 48, 49, 50],
@@ III Summary of results
Figure 2: Expected phase diagrams for (a) the Kapit-Mueller model and (b) the checkerboard model with the pinning potential $V_{p}\neq 0$. $\phi$ is the applied magnetic field and $U$ is the strength of the nearest-neighbor interaction. The strength of the pinning potential is chosen so that the created quasiparticles are well localized. It is assumed that effects of interactions between the localized quasiparticles are negligible in a thermodynamically large system.
Before presenting detailed calculation results, we first summarize the key results here. quasiparticles can be created by the magnetic field $\phi$, and they can be localized by an additional pinning potential $V_{p}$ corresponding to disorder. We show expected phase diagrams of the two models with the pinning potential $V_{p}\neq 0$ in Fig. 2 based on the exact diagonalization and the quantum geometry under magnetic fields. We have used different values of $V_{p}$ in the exact diagonalization and details of the phase diagram could depend on the strength of $V_{p}$, although it is qualitatively unchanged. Generally, a magnetic field increases or decreases the number of (nearly degenerate) single-particle states $N_{\text{orb}}$ according to the Streda’s formula [53, 54, 55, 56], $N_{\text{orb}}\to N_{\text{orb}}+C_{s}N_{\phi}$, where $C_{\text{s}}$ is the Chern number of the band considered and $N_{\phi}$ is the number of the added flux for the entire system. If $N_{\text{orb}}$ increases (decreases) under a magnetic field, the filling $\nu=N/N_{\text{orb}}$ decreases (increases), which can lead to quasihole (quasielectron) creation. In the Kapit-Mueller model, since the Chern number of the lowest energy band is $+1$, adding a positive magnetic field induces quasihole creation, while a negative magnetic field leads to quasielectron creation. In the checkerboard model, the Chern number of the lowest energy band is $-1$, and a positive (negative) magnetic field corresponds to quasielectron (quasihole) creation. The quasiholes can be localized by a pinning potential $V_{p}>0$ while the quasielectrons are by $V_{p}<0$ in our notation. Within the exact diagonalization for limited system sizes in the present study, the number of the created quasiparticles is $N_{\text{qp}}=1$ in each model. In a thermodynamically large system, the number of the flux quanta is $|N_{\phi}|\gg 1$ for a fixed magnetic field $\phi$ and there will be $|N_{\phi}|$ quasiparticles which are localized by a pinning potential. In Fig. 2, we have assumed that the quasiparticles are well localized and effects of interactions between the pinned quasiparticles are negligible.
As seen in Fig. 2(a) for the Kapit-Mueller model, the FCI state is stable under the magnetic field irrespective of its direction. More precisely, except for very small interactions, quasiparticles are stably created by the magnetic field and then localized by the pinning potential. Consequently, the fractionally quantized Hall conductance is maintained over a wide range of parameters, which means that the FCI phase is stable. The phase diagram is almost symmetric for $+\phi$ and $-\phi$, and thus the creations of quasiholes ($\phi>0$) and quasielectrons ($\phi<0$) are both stable. This is a reasonable result, since the Kapit-Mueller model is a lattice model which well mimics the lowest Landau level in the continuum system and basic properties are similar to the corresponding FQHE.
On the other hand, in the checkerboard model, the quasielectron creation under $\phi>0$ is stable except for small interactions, while the quasihole creation for $\phi<0$ is unstable even for moderately strong interactions. In the latter case, the FCI state is destabilized by the magnetic field and a non-FCI state is realized. The non-FCI state at very small interactions is smoothly connected to the non-interacting Fermi liquid at $U=0$. However, it is difficult to fully clarify nature of the non-FCI state within the present calculations, and exact identification of the non-FCI state is left for a future study. We stress that the asymmetry with respect to ±$\pm\phi$ is a characteristic feature of a lattice system which is absent in the standard FQHE in the continuum model.
@@ IV.1 Characterization of FCI
In this section, we study the FCI with use of the exact diagonalization. First, we briefly explain how to characterize an FCI state within the exact diagonalization. To examine stability of the FCI states against magnetic fields, we apply one flux quantum $N_{\phi}=\pm 1$ uniformly over the entire system and discuss quasiparticles induced by the magnetic field based on the “counting-rule” on the energy spectra [1, 57, 58]. Previous works examined quasihole and quasielectron creations in FCIs at $\phi=0$ based on the counting-rule, when the number of unit cells is increased or decreased [41, 42, 43]. When the system size is changed with a fixed number $N$ of the fermions, the number of the nearly degenerate single-particle states $N_{\text{orb}}$ and the filling $\nu=N/N_{\text{orb}}$ varies, while the single-particle band structure itself is unchanged. Instead of varying the system size, a change in the number of the single-particle orbits can be induced by the magnetic field according to the Streda’s formula, $N_{\text{orb}}\to N_{\text{orb}}+C_{s}N_{\phi}$. In contrast to the system-size induced change of $N_{\text{orb}}$, the system does not keep the (magnetic) translation symmetry under the additional one flux quantum $N_{\phi}=\pm 1$ [49]. However, one can still regard the nearly degenerate single-particle orbits as a nearly flat band and discuss the stability of the FCI in a manner similar to those in the previous studies at $\phi=0$.
Since the single-particle Chern number is $|C_{s}|=1$ in the present models, a single flux quantum leads to the change $N_{\text{orb}}\to N_{\text{orb}}\pm 1$. This gives rise to creation of one quasihole or quasielectron with the fractional charge $|e^{*}|=1/3$. In case of the quasihole creation, the ground state degeneracy #$\#_{\text{qh}}$ under the magnetic field obeys the counting rule given by  [1]
@@
where $N_{\text{orb}}$ is the number of the low energy single-particle states in the presence of the magnetic field. For quasielectron creation, no simple formula for the ground state degeneracy #$\#_{\text{qe}}$ is known, but we can verify the counting rule of the quasielectrons by referring to energy spectra of the corresponding FQHE states. More precisely, we calculate #$\#_{\text{qe}}$ for the LLL in the electron gas model and also for fermions in the Kapit-Mueller model with the Poisson summation rule which is a lattice realization of the LLL [23, 46]. We numerically find that #$\#_{\text{qe}}=11$ for $(N=4,N_{\text{orb}}=11)$, #$\#_{\text{qe}}=14$ for $(N=5,N_{\text{orb}}=14)$, and #$\#_{\text{qe}}=17$ for $(N=6,N_{\text{orb}}=17)$.
The quasiparticles created by the uniform magnetic field are extended over the entire system. They can be spatially localized in the presence of an additional pinning potential $V_{p}$ which couples with the quasiparticle charge [39, 40]. To this end, we first perform the exact diagonalization to obtain low energy states |$\ket{\Psi_{1}},\ket{\Psi_{2}},\cdots,\ket{\Psi_{N_{\text{cut}}}}$ without the pinning potential [41, 42], where $N_{\text{cut}}=\#_{\text{qh}}$ or #$\#_{\text{qe}}$ is a cut-off for the low energy states. Then, we take $V_{p}$ into account within this low energy sector by diagonalizing the projected Hamiltonian with the pinning potential,
@@
where $H$ is either $H_{\text{KM}}$ or $H_{\text{CB}}$. $P$ is the projection onto the low energy states {|$\{\ket{\Psi_{n}}\}_{n=1}^{N_{\text{cut}}}$ and $j_{0}$ is the position of the pinning potential. The projection enables us to efficiently focus on the low energy sector and also greatly reduces computational costs for the analysis of the pinning potential [41, 42]. When the quasiparticle created by $\phi\neq 0$ is localized by the pinning potential, it does not contribute to the Hall conductance. Correspondingly, the ground state degeneracy of $\tilde{H}$ will be three fold for the present $\nu=1/3$ system. In this sense, the ground states of $\tilde{H}$ belong to the same FCI phase of $H$ at $\phi=0$. However, the quasiparticle creation can be unstable and a magnetic field may destabilize the underlying FCI state. In this case, the counting rule does not hold and the ground states of $\tilde{H}$ are no longer FCIs.
@@
The variance is exponentially small if the ground states are FCIs, while it can be large for gapless states. In this way, we can identify FCI states within the exact diagonalization for small system sizes. It turns out that, even if the ground state of $\tilde{H}$ is an FCI, $\mathcal{V}[\mathcal{C}]$ diverges when we use $N_{\text{cut}}<\#_{\text{qh}},\#_{\text{qe}}$. When $N_{\text{cut}}=\#_{\text{qh}},\#_{\text{qe}}$, the variance $\mathcal{V}[\mathcal{C}]$ is well suppressed in the FCI phase. This is because it is necessary to include all the states |$\ket{\Psi_{1}},\cdots,\ket{\Psi_{\#{\rm qh(qe)}}}$ with the spatially extended quasiparticle at $V_{p}=0$ for its localization under $V_{p}\neq 0$.
In this study, we consider the systems with the number of the fermions $N=4\sim 6$. This is relatively small compared to those in the previous studies where quasiparticles are created by the increase of the system sizes [1, 41, 42, 43]. Technically, one flux quantum breaks the (magnetic) translation symmetry and the size of the Hilbert space cannot be reduced by the translation symmetry [49]. However, it turns that we can discuss physical properties of the systems with $N=4\sim 6$ by carefully paying attention to finite size effects as will be discussed in the following sections. Furthermore, we have confirmed that quasiparticles are stably created by the increase of the system sizes for $N=4,5$ in consistent with the previous studies [41, 42, 43]. See also Appendix A. This implies that finite size effects are small for these system sizes.